OpenStudy (ny,ny):

Solve each system using substitution. Check your solution. h=6g-4 h=-2g+28

OpenStudy (anonymous):

2.h= 6g-4...(i) h= -2g+28...(ii) 6g-4=-2g+28 +2g +2g 8g-4=28 8g=32 g=32/8 if you not - sign solve for h h=6g-4

OpenStudy (ny,ny):

g=4.

OpenStudy (anonymous):

right

OpenStudy (ny,ny):

so how do I solve for h?

OpenStudy (anonymous):

im not sure.. thats the only part i no;) sry..

OpenStudy (ny,ny):

alright. thanks anyway =)

OpenStudy (anonymous):

np=D

OpenStudy (anonymous):

medal me?;)

OpenStudy (ny,ny):

i fanned you. i want to give medal to person who solves all 3 parts of the problem. only fair that way. i appreciate your help though. :D

OpenStudy (anonymous):

since you know g now, plug it into one of the original equations and solve for h

OpenStudy (anonymous):

h=6(4)-4 h=20

OpenStudy (anonymous):

To check your solutions, plug both of your answers into one of the original equations and if its true, you were right 20=6(4)-4 20=20 So the answers are right. h=20 g=4

OpenStudy (ny,ny):

well yeah, i know i could just plug it in, but my teacher taught me another way. i had to replace and all this other stuff. do you know that way?

OpenStudy (anonymous):

Well to solve for h that way, do the same thing you did for g.

OpenStudy (ny,ny):

the first thing i did to solve for g was 6g-4=-2g+28 how do i do that to solve for h?

OpenStudy (anonymous):

First solve one of the equations for g. h=6g-4 h+4=6g (h+4)/6=g now replace g in your other equation with (h+4)/6 h=-2((h+4)/6)+28 and solve for h.

OpenStudy (ny,ny):

h=20. ok. well i was looking through my notes because the other way that i had asked for you to help me with was unfamiliar with me. so many numbers. i just found out that the first way you had showed me was the way our teacher would have solved the problem. im sorry i made you wait so long and that i made you write all this stuff. Thanks for all your help. I understand what to do now.

OpenStudy (anonymous):

You're welcome.